### v1

@@ 1. Introduction
While RTA theoretically centers the researcher’s positionality and the evolution of meaning (Probst and Berenson, 2014), achieving this rigor in practice remains difficult. In current workflows, critical reflexive notes are often disconnected from the data; the evolving nature of codes is rarely documented in a provenance-driven manner; and collaborator positionalities remain invisible when interpretations diverge (Singh et al., 2025).
Compounding this challenge, the rapid growth of tool support for qualitative data analysis (QDA) has largely pulled in the opposite direction. While commercial tools like NVivo and ATLAS.ti facilitate management and prioritize quantitative insights, recent research systems leveraging Large Language Models (LLMs) have prioritized automation, scalability, and convergence. New tools often act as “auto-coders” to reduce human effort (Gao et al., 2023; Ye et al., 2025) or mediate discussions to reach rapid consensus (Gao et al., 2024).
@@
We address this misalignment with Reflexis, a collaborative workspace designed to embed reflexivity, analytic history, and positionality directly into the analysis loop. Crucially, our goal is not to automate the coding process for speed, nor to enforce inter-coder reliability. Instead, we aim to scaffold methodological rigor by designing for deliberation (Ma et al., 2025). Reflexis supports this through three principal mechanisms: (1) In-situ Reflexivity: Instead of separating memos from coding, in-situ reflective prompts link researchers’ evolving perspectives to specific data excerpts without disrupting their analytical flow. (2) Transparent Provenance: To counter the opacity of evolving concepts, a transparent analysis history records the full provenance of every code’s evolution (creations, splits, merges), while a code drift alert surfaces potential inconsistencies for human review. (3) Principled Disagreement: When interpretations diverge, positionality-aware collaboration and discussion focus tools surface these differences as opportunities for deeper theorization rather than errors to be fixed.
Consistent with RTA’s interpretivist lens, all AI assistance in Reflexis (e.g., detecting drift) is advisory and optional, ensuring the system supports, and not supplants human interpretation. Guided by this stance, we investigate the effectiveness of Reflexis in support of collaborative RTA via the following research questions:
@@ 3. RQ3
    (Scaffolding Collaborative Interpretation): How do positionality-aware scaffolds help research pairs identify and frame interpretive differences during synchronous collaborative workflow?
@@
This paper contributes: (1) A conceptual framework for Design for Deliberation in Qualitative Analysis, challenging the prevailing focus on automation and convergence by articulating how AI can be used to scaffold friction, reflexivity, and interpretive diversity. (2) Reflexis, a collaborative system that operationalizes this deliberative approach within RTA. It introduces novel interaction mechanisms that transform implicit mental processes of reflection into explicit, tangible artifacts. (3) Empirical insights from an evaluation study demonstrating how deliberative designs shape analytical practice. Our work illustrates the value of positioning AI as a partner in rigor and deliberation. Our findings indicate that this approach helps teams reframe disagreements as opportunities for deeper qualitative insights rather than obstacles to efficiency.
@@ 2.1. Reflexivity in Qualitative Analysis
Despite its theoretical clarity, significant challenges arise in the application and evaluation of RTA within the field of human-computer interaction (HCI). A primary issue is the inconsistent reporting of reflexive practices. A recent review found that most HCI papers claiming an RTA approach failed to adequately describe the researchers’ positions, leaving readers to infer the interpretive lens through which the data was analyzed (McDonald et al., 2019). This lack of methodological transparency is a critical problem for qualitative research, where understanding the perspective of the analyst is paramount (Moravcsik, 2014). A related challenge is positivism creep, where interpretivist work is judged by criteria from a positivist/quantitative research paradigm (Soden et al., 2024; Bowman et al., 2023). Reviewers may request justifications for participant counts, demographic tables, or inter-coder reliability scores, which are metrics that are misaligned with interpretive validity (Braun and Clarke, 2016; Noble and Smith, 2015). This pressure can result in “performative” positionality statements that do little to illuminate the research process (Singh et al., 2025). This environment places researchers in a difficult position, as they must advocate for the value of interpretation while navigating demands for positivist “rigor” (Watkins, 2017), a challenge that is intensified in collaborative team settings where managing reflexivity is inherently more complex (Barry et al., 1999; Probst and Berenson, 2014).
@@ 2.2. The Role of AI in Qualitative Analysis
This rapid pace of tool development, however, often leaves critical tensions for reflexive methods like RTA unaddressed. A primary concern is that AI assistance can distance researchers from their data, inhibiting the deep immersion and critical questioning, that is foundational to the interpretive process (Jiang et al., 2021). Using AI to automate interpretation also risks supplanting human agency and the critical, interpretive lens, a significant problem for reflexive analysis that often has a social justice orientation (Braun and Clarke, 2019). Furthermore, introducing opaque computational systems complicates assessments of methodological rigor and creates challenges for establishing appropriate trust and reliance on automation (Lee and See, 2004; McDonald et al., 2019). In response, scholars have proposed that instead of seeking consensus, AI tools should be designed to highlight ambiguity and disagreement as sites for reflexive engagement (Jiang et al., 2021; Chen et al., 2018). Others have put forth design principles for ethical LLM integration, calling for tools that enable intentional use in specific subtasks, offer transparency, and account for researcher positionality (Schroeder et al., 2025). We build on this work by designing and evaluating a system that operationalizes these principles, focusing on an LLM-powered tool that centers researcher reflexivity and supports the communication of analytic rigor.
@@ 2.3. Collaboration, Reliability and Provenance in Qualitative Analysis
Collaborative qualitative analysis involves complex coordination, requiring strategies to manage teamwork and build shared understanding, often across physical distances (Cornish et al., 2014; Hall et al., 2005; Richards and Hemphill, 2018; Olson and Olson, 2000). A key site of tension within this process is the concept of reliability, particularly the use of inter-coder reliability (ICR) metrics. However, the role and appropriateness of ICR for validating qualitative conclusions (Kurasaki, 2000), have become a subject of considerable debate (O’Connor and Joffe, 2020). Many scholars argue that its positivist roots are fundamentally incompatible with the interpretive goals of “Big-Q” methods like RTA (Kidder and Fine, 1987), a sentiment reflected in the low reporting rates of ICR in HCI and CSCW literature (O’Connor and Joffe, 2020; McDonald et al., 2019; Díaz et al., 2023). However, an alternative perspective reframes disagreement not as the failure to attain reliability, but as an analytically productive event that signals opportunities for deeper insight (Zade et al., 2018). From this view, the purpose and utility of reliability metrics should be re-envisioned. Instead of serving as a definitive measure of interpretive agreement, they can be used as a diagnostic tool to spark reflexive conversations about why interpretations differ (O’Connor and Joffe, 2020).
In addition, a key component of managing collaboration and ensuring methodological rigor is maintaining a transparent history of how analytical concepts evolve. Prior HCI and information visualization communities have explored analytical provenance (North et al., 2011; Ragan et al., 2015). While early HCI work visualized collaborative document evolution (Viégas et al., 2004; Wang et al., 2015), visualization research distinguishes between capturing low-level user actions (interaction provenance) (Huang and Nobre, 2025; Cutler et al., 2020) and higher-level reasoning processes (insight provenance) (Gotz and Zhou, 2009). This distinction is critical for qualitative analysis, as interaction logs alone cannot capture the rationale behind interpretive shifts. Research in visual sensemaking suggests that without explicit scaffolds to track these shifts, which are termed “reframing” in the Data-Frame theory (Klein et al., 2007), analysts often struggle to recall the reasoning behind their earlier decisions (Lipford et al., 2010).
@@ 3. Formative Study
We surveyed Prolific, targeting postgraduate researchers to understand current reflexive practices and collaborative challenges in qualitative analysis. The 30-minute preliminary survey examined how researchers engage in reflexivity during coding, track code evolution over time, and navigate differences in collaborative interpretation. We screened for familiarity with TA and prior qualitative analysis experience, retaining 55 valid responses. Participants were compensated at \$16/hour (approximately \$8 for the 30-minute survey). Details for survey questions, Prolific screening questions and survey completion statistics can be found in Appendix A.1.
Survey responses were analyzed using descriptive statistics for Likert-scale questions and thematic coding for open-ended responses. The first author conducted initial analysis, with findings reviewed by the second author to ensure reliability. While the survey confirmed that researchers value reflexive practices and recognize their importance for analytical rigour, responses revealed significant implementation challenges. For example, one respondent noted: “my own assumptions can get in the way, I have to keep checking myself” ($S_{44}$). These responses suggested that understanding the subtle workflow barriers preventing systematic reflexivity required deeper exploration than survey methods could provide.
@@
To gain a more in-depth understanding of these implementation challenges, we conducted one-hour semi-structured interviews with three experienced RTA practitioners (2 PhD students, 1 university professor). Participants were invited through the authors’ professional and academic networks. We employed purposive sampling, a non-probability technique where participants are deliberately selected based on specific characteristics relevant to the research question (Palinkas et al., 2015). This approach was appropriate here because it ensured interviewees possessed the deep RTA expertise needed to provide nuanced insights into reflexive practices that may not be evident from survey responses alone. This sampling strategy allowed us to directly verify that all practitioners met our criteria for adequate RTA experience. [^3] All three interview participants have completed at least two formal qualitative studies as part of their research portfolio, with at least one using reflexive thematic analysis. Additionally, two of the graduate students have been formally trained in reflexive thematic analysis through a graduate-level course. The university professor has extensive experience publishing research work in reputable peer-reviewed conferences/journals in their research domain, using reflexive thematic analysis as a method. All three participants have different research focuses within different sub-disciplines of HCI.
All interviews were conducted virtually via Zoom and facilitated by the first author. Participants received a \$20 for their time. Interviews were recorded for analysis. Recorded interviews began with participants describing a recent workflow to ground the discussion in concrete practices. We then explored their approaches to reflexivity, code evolution, and collaborative interpretation. Finally, participants evaluated early system concepts and wireframes. The complete interview script is documented in Appendix A.6.1.
@@
3.1. Challenges Identified from Existing Behaviours
@@
Table 1. Mapping from user challenges and behaviours to design goals and system features.
Our survey and interview data reveal that researchers actively engage in reflexive practices but struggle with tools that do not support the non-linear nature of RTA. Table A.1 maps these observed behaviours to design goals and system features.
@@ 3.1.1. Challenge 1: Inconsistent Reflection Support
Our survey reveals that qualitative researchers desire deep reflexivity, but implementation is sporadic. While respondents ($40/55$) “often” or “always” consider how their experience shapes interpretation, they lack integrated tools to capture it. Researchers rated “asking myself guiding questions” as highly helpful, yet currently rely on disconnected methods like “writing a separate research journal” ($32/55$). As $S_{44}$ noted, “my own assumptions can get in the way,” yet without integrated support, checking these assumptions becomes a disruption.
@@
To address this fragmentation, participants suggested distinct alternatives. $I_{1}$ advocated for a “spatial whiteboarding” interface similar to LiquidText (Tashman and Edwards, 2011) to visually map connections, while $I_{3}$ suggested “stage-based checklists” enforced before and after coding sessions. However, we reasoned that spatial tools often separate analysis from the raw text, and stage-based checks miss the immediate moment of interpretive decision-making. Consequently, we prioritized in-the-moment reflexivity support (DG1) to keep the raw text visible. We integrated $I_{2}$’s suggestion for a user-initiated reflection, which allows researchers to engage in reflection on their own pace, with self-selected reflection areas, rather than time-based or progress-based triggers.
@@ 3.1.2. Challenge 2: Untracked Code Evolution
Although researchers recognize the importance of tracking how codes shift, current approaches are manual and reactive. To manage code drift, researchers update definitions ($37/55$) or split codes ($31/55$), often relying on memory ($29/55$) to maintain consistency. $S_{17}$ described maintaining a “living codebook,” while $S_{3}$ used a “separate tab for the initial code book” to duplicate data. These ad-hoc workarounds often lead to incomplete documentation, with $I_{1}$ noting the desire for detection only after “a significant amount of coding has already been done.”
@@
Participants emphasized the need to control how code definitions change. $I_{2}$ suggested implementing hierarchical role-based visibility, where a “project lead” blinds junior researchers to specific codes to ensure independent thought. We addressed this by incorporating flexible visibility controls that allow users to selectively display or hide analysis progress. However, we reasoned that while hiding information can reduce bias, maintaining rigor requires the ability to audit how those definitions shifted over time. Therefore, we prioritized transparent provenance (DG2) as the underlying framework, adopting $I_{1}$’s suggestion to treat code history as an exportable artifact useful for open science and pedagogy.
@@ 3.1.3. Challenge 3: Inaccessible Collaboration Perspectives
Accessing and integrating collaborator positionalities remains a persistent friction point. While respondents value understanding a partner’s “disciplinary perspective” ($35/55$) and “relevant lived experiences” ($37/55$), this information is rarely integrated into the coding interface. $S_{34}$ highlighted how “communication breakdowns” occur when trying to ensure consistency while respecting diverse perspectives. Currently, teams rely on informal discussions to surface these lenses, which $I_{1}$ noted leads to inconsistent awareness of “why” interpretations differ.
@@
Participants proposed specific reflexive interventions, such as “fairness audits” ($I_{3}$) or “dominant narrative” critiques ($I_{1}$), where researchers explicitly examine how their findings fit within or resist the prevailing discourse of the field. We reasoned that hard-coding these specific theoretical frameworks might limit the tool’s applicability across different RTA contexts. Instead, we focused on the structural mechanism of “double coding” proposed by $I_{3}$, which treats conflicting interpretations as valid and parallel data points rather than errors to be resolved. We aimed to surface principled disagreement (DG3) by displaying these divergent codes side-by-side with researcher profiles. This facilitates the “why not” analysis ($I_{3}$) where users consider why they missed a perspective seen by their partner.

### v2

@@ 1. Introduction
While RTA theoretically centers the researcher’s positionality and the evolution of meaning (Probst and Berenson, 2014), achieving this rigor in practice remains difficult. In current workflows, self-reflexive notes are often disconnected from the data; the evolving nature of codes is rarely documented with provenance; and collaborator positionalities remain invisible when interpretations diverge (Singh et al., 2025).
Compounding this challenge is the rapid growth of tool support for qualitative data analysis (QDA) that has largely pulled in the opposite direction. While commercial tools like NVivo and ATLAS.ti facilitate management and prioritize quantitative insights, recent research systems leveraging Large Language Models (LLMs) have prioritized automation, scalability, and convergence. New tools often act as “auto-coders” to reduce human effort (Gao et al., 2023; Ye et al., 2025) or mediate discussions to reach rapid consensus (Gao et al., 2024).
@@
We address this misalignment with Reflexis, a collaborative workspace designed to embed reflexivity, analytic history, and positionality directly into the analysis loop. Crucially, our goal is not to automate the coding process for speed, nor to enforce inter-coder reliability. Instead, we aim to scaffold methodological rigor by designing for deliberation (Ma et al., 2025). Reflexis supports this through three principal mechanisms: (1) In-situ Reflexivity: Instead of separating memos from coding, in-situ reflective prompts link researchers’ evolving perspectives to specific data excerpts without disrupting their analytical flow. (2) Transparent Provenance: To counter the opacity of evolving concepts, a transparent analysis history records the full provenance of every code’s evolution (creations, splits, merges), while a Code Drift Alert surfaces potential inconsistencies for human review. (3) Principled Disagreement: When interpretations diverge, positionality-aware collaboration and Discussion Focus surface these differences as opportunities for deeper theorization rather than errors to be fixed.
Consistent with RTA’s interpretivist lens, all AI assistance in Reflexis (e.g., detecting drift) is advisory and optional, ensuring the system supports, rather than supplanting, human interpretation. Guided by this stance, we investigate the effectiveness of Reflexis in support of collaborative RTA via the following research questions:
@@ 3. RQ3
    (Scaffolding Collaborative Interpretation): How do positionality-aware scaffolds help research pairs identify and frame interpretive differences during synchronous collaborative work?
@@
This paper contributes: (1) A conceptual framework for Design for Deliberation in Qualitative Analysis, challenging the prevailing focus on automation and convergence by articulating how AI can be used to scaffold friction, reflexivity, and interpretive diversity. (2) Reflexis, a collaborative system that operationalizes this approach within RTA. It introduces novel interaction mechanisms that transform implicit mental processes of reflection into explicit, tangible artifacts. (3) Empirical insights from an evaluation study demonstrating how deliberative designs shape analytical practice. Our work illustrates the value of positioning AI as a partner in rigor and deliberation. Our findings suggest that this approach helps teams reframe disagreements as opportunities for deeper qualitative insights.
@@ 2.1. Reflexivity in Qualitative Analysis
Despite its theoretical clarity, significant challenges arise in the application and evaluation of RTA within the field of human-computer interaction (HCI). A primary issue is the inconsistent reporting of reflexive practices. A recent review has found that most HCI papers claiming an RTA approach failed to adequately describe the researchers’ positions, leaving readers to infer the interpretive lens through which the data was analyzed (McDonald et al., 2019). This lack of methodological transparency is a critical problem for qualitative research, where understanding the perspective of the analyst is paramount (Moravcsik, 2014). A related challenge is positivism creep, where interpretivist work is judged by criteria from a positivist/quantitative research paradigm (Soden et al., 2024; Bowman et al., 2023). Reviewers may request justifications for participant counts, demographic tables, or inter-coder reliability scores, metrics that are misaligned with interpretive validity (Braun and Clarke, 2016; Noble and Smith, 2015). This pressure can result in “performative” positionality statements that do little to illuminate the research process (Singh et al., 2025). This environment places researchers in a difficult position, as they must advocate for the value of interpretation while navigating demands for positivist “rigor” (Watkins, 2017), a challenge that is intensified in collaborative team settings where managing reflexivity is inherently more complex (Barry et al., 1999; Probst and Berenson, 2014).
@@ 2.2. The Role of AI in Qualitative Analysis
This rapid pace of tool development, however, often leaves critical tensions for reflexive methods like RTA unaddressed. A primary concern is that AI assistance can distance researchers from their data, inhibiting the deep immersion and critical questioning that is foundational to the interpretive process (Jiang et al., 2021). Using AI to automate interpretation also risks supplanting human agency and the critical, interpretive lens, a significant problem for reflexive analysis that often has a social justice orientation (Braun and Clarke, 2019). Furthermore, introducing opaque computational systems complicates assessments of methodological rigor and creates challenges for establishing appropriate trust and reliance on automation (Lee and See, 2004; McDonald et al., 2019). In response, scholars have proposed that instead of seeking consensus, AI tools should be designed to highlight ambiguity and disagreement as sites for reflexive engagement (Jiang et al., 2021; Chen et al., 2018). Others have put forth design principles for ethical LLM integration, calling for tools that enable intentional use in specific subtasks, offer transparency, and account for researcher positionality (Schroeder et al., 2025). We build on this work by designing and evaluating a system that operationalizes these principles, focusing on an LLM-powered tool that centers researcher reflexivity and supports the communication of analytic rigor.
@@ 2.3. Collaboration, Reliability and Provenance in Qualitative Analysis
Collaborative qualitative analysis involves complex coordination, requiring strategies to manage teamwork and build shared understanding, often across physical distances (Cornish et al., 2014; Hall et al., 2005; Richards and Hemphill, 2018; Olson and Olson, 2000). A key site of tension within this process is the concept of reliability, particularly the use of inter-coder reliability (ICR) metrics. The role and appropriateness of ICR for validating qualitative conclusions (Kurasaki, 2000) has become a subject of considerable debate (O’Connor and Joffe, 2020). Many scholars argue that its positivist roots are fundamentally incompatible with the interpretive goals of “Big-Q” methods like RTA (Kidder and Fine, 1987), a sentiment reflected in the low reporting rates of ICR in HCI and CSCW literature (O’Connor and Joffe, 2020; McDonald et al., 2019; Díaz et al., 2023). However, an alternative perspective reframes disagreement not as the failure to attain reliability, but as an analytically productive event that signals opportunities for deeper insight (Zade et al., 2018). From this view, the purpose and utility of reliability metrics should be re-envisioned. Instead of serving as a definitive measure of interpretive agreement, they can be used as a diagnostic tool to spark reflexive conversations about why interpretations differ (O’Connor and Joffe, 2020).
In addition, a key component of managing collaboration and ensuring methodological rigor is maintaining a transparent history of how analytical concepts evolve. Prior HCI and information visualization communities have explored analytical provenance (North et al., 2011; Ragan et al., 2015). While early HCI work visualized collaborative document evolution (Viégas et al., 2004; Wang et al., 2015), visualization research distinguishes between capturing low-level user actions (interaction provenance) (Huang and Nobre, 2025; Cutler et al., 2020) and higher-level reasoning processes (insight provenance) (Gotz and Zhou, 2009). This distinction is critical for qualitative analysis, as interaction logs alone cannot capture the rationale behind interpretive shifts. Research in visual sensemaking suggests that without explicit scaffolds to track these shifts, which are termed “reframing” in Data-Frame theory (Klein et al., 2007), analysts often struggle to recall the reasoning behind their earlier decisions (Lipford et al., 2010).
@@ 3. Formative Study
We conducted our survey on Prolific, targeting postgraduate researchers to understand current reflexive practices and collaborative challenges in qualitative analysis. The 30-minute preliminary survey examined how researchers engage in reflexivity during coding, track code evolution over time, and navigate differences in collaborative interpretation. We screened for familiarity with TA and prior qualitative analysis experience, retaining 55 valid responses. Participants were compensated at \$16/hour (approximately \$8 for the 30-minute survey). The protocol, consent, and compensation were approved by our university IRB office (protocol # 49033). Details about survey questions, Prolific screening questions, and survey completion statistics can be found in Appendix A.1.
Survey responses were analyzed using descriptive statistics for Likert-scale questions and thematic coding for open-ended responses. The first author conducted the initial analysis, and the second author reviewed the findings to ensure reliability. While the survey confirmed that researchers value reflexive practices and recognize their importance for analytical rigor, responses revealed significant implementation challenges. For example, one respondent noted: “my own assumptions can get in the way, I have to keep checking myself” ($S_{44}$). These responses suggested that understanding the subtle workflow barriers preventing systematic reflexivity required deeper exploration than survey methods could provide.
@@
To gain a more in-depth understanding of these implementation challenges, we conducted one-hour semi-structured interviews with three experienced RTA practitioners (2 PhD students, 1 university professor). Participants were invited through the authors’ professional and academic networks. We employed purposive sampling, a non-probabilistic technique where participants are deliberately selected based on specific characteristics relevant to the research question (Palinkas et al., 2015). We deemed this approach appropriate because it ensured interviewees possessed the deep RTA expertise needed to provide nuanced insights into reflexive practices that may not be evident from survey responses alone. This sampling strategy allowed us to directly verify that all practitioners met our criteria for adequate RTA experience. [^3] All three interview participants have completed at least two formal qualitative studies as part of their research portfolio, with at least one using reflexive thematic analysis. Additionally, two of the graduate students have been formally trained in reflexive thematic analysis through a graduate-level course. The university professor has extensive experience publishing research work in reputable peer-reviewed conferences/journals in their research domain, using reflexive thematic analysis as a method. All three participants have different research foci within different sub-disciplines of HCI.
All interviews were conducted virtually via Zoom and facilitated by the first author. Participants received \$20 for their time. Interviews were recorded for analysis. Recorded interviews began with participants describing a recent workflow to ground the discussion in concrete practices. We then explored their approaches to reflexivity, code evolution, and collaborative interpretation. Finally, participants evaluated early system concepts and wireframes. The complete interview script is documented in Appendix A.6.1.
@@
3.1. Challenges Identified from Existing Behaviors
@@
Table 1. Mapping from user challenges and behaviors to design goals and system features.
Our survey and interview data reveal that researchers actively engage in reflexive practices but struggle with tools that do not support the non-linear nature of RTA. Table A.1 maps these observed behaviors to design goals and system features.
@@ 3.1.1. Challenge 1: Inconsistent Reflection Support
Our survey revealed that qualitative researchers desire deep reflexivity, but implementation is sporadic. While respondents ($40/55$) “often” or “always” considered how their experience shaped interpretation, they lacked integrated tools to capture it. Researchers rated “asking myself guiding questions” as highly helpful, yet currently rely on disconnected methods like “writing a separate research journal” ($32/55$). As $S_{44}$ noted, “my own assumptions can get in the way,” yet without integrated support, checking these assumptions became a disruption.
@@
To address this fragmentation, participants suggested distinct alternatives. $I_{1}$ advocated for a “spatial whiteboarding” interface similar to LiquidText (Tashman and Edwards, 2011) to visually map connections, while $I_{3}$ suggested “stage-based checklists” enforced before and after coding sessions. However, we reasoned that spatial tools often separated analysis from the raw text, and stage-based checks missed the immediate moment of interpretive decision-making. Consequently, we prioritized in-the-moment reflexivity support (DG1) to keep the raw text visible. We integrated $I_{2}$’s suggestion for a user-initiated reflection, which allowed researchers to engage in reflection at their own pace with self-selected reflection areas, rather than time-based or progress-based triggers.
@@ 3.1.2. Challenge 2: Untracked Code Evolution
Although researchers recognized the importance of tracking how codes shift, current approaches were manual and reactive. To manage code drift, researchers updated definitions ($37/55$) or split codes ($31/55$), often relying on memory ($29/55$) to maintain consistency. $S_{17}$ described maintaining a “living codebook,” while $S_{3}$ used a “separate tab for the initial code book” to duplicate data.
@@
Participants emphasized the need to control how code definitions change. $I_{2}$ suggested implementing hierarchical role-based visibility, where a “project lead” blinds junior researchers to specific codes to ensure independent thought. We addressed this by incorporating flexible visibility controls that allow users to selectively display or hide analysis progress. However, we reasoned that while hiding information could reduce bias, maintaining rigor required the ability to audit how those definitions shifted over time. Therefore, we prioritized transparent provenance (DG2) as the underlying framework, adopting $I_{1}$’s suggestion to treat code history as an exportable artifact useful for open science and pedagogy.
@@ 3.1.3. Challenge 3: Inaccessible Collaboration Perspectives
Accessing and integrating collaborator positionalities remained a persistent friction point. While respondents valued understanding a partner’s “disciplinary perspective” ($35/55$) and “relevant lived experiences” ($37/55$), this information was rarely integrated into the coding interface. $S_{34}$ highlighted how “communication breakdowns” occur when trying to ensure consistency while respecting diverse perspectives. Currently, teams rely on informal discussions to surface these lenses, which $I_{1}$ noted led to inconsistent awareness of “why” interpretations differ.
@@
Participants proposed specific reflexive interventions, such as “fairness audits” ($I_{3}$) or “dominant narrative” critiques ($I_{1}$), where researchers explicitly examined how their findings fit within or resist the prevailing discourse of the field. We reasoned that hard-coding these specific theoretical frameworks might limit the tool’s applicability across different RTA contexts. Instead, we focused on the structural mechanism of “double coding” proposed by $I_{3}$, which treated conflicting interpretations as valid and parallel data points rather than errors to be resolved. We aimed to surface principled disagreement (DG3) by displaying these divergent codes side-by-side with researcher profiles. This facilitated the “why not” analysis ($I_{3}$), in which users considered why they missed a perspective held by their partner.